OpenStudy (anonymous):

Brooke collects porcelain dolls and teddy bears. Over the years, the number of porcelain dolls she collected is 14 less than twice the number of teddy bears. If she has a total of 52 dolls and bears, how many teddy bears does Brooke have? help plz. will give medal

OpenStudy (cheesecakekitten):

Too many

OpenStudy (anonymous):

lol plz help

OpenStudy (anonymous):

52/2=26 26-14=12 She has 12 porcelain dolls. 52-12=40 teddy bears

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

You're welcome! Anytime. :) @flvsisstupid